Given a firms demand function Q-90+2P=0 and it's average cost function AC=Q²-8Q+57+2/Q. Find the level of output which maximises marginal cost.




1
Expert's answer
2021-05-25T11:14:28-0400

Average cost function"=AC=Q^2-8Q+57+\frac{2}{Q}"

"TC=AC\times\ Q=Q^3-8Q^2+57Q+2"

"MC=\frac{dTC}{dQ}=3Q^2-16Q+57"

During the first order condition, "MC=0"

Therefore, "3Q^2-16Q+57=0"

Solving for "Q" using the quadratic equation we get "Q=-4.68\ or \ 36.68"

36.68 will maximize Marginal Cost.